40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 17     Entry time: Wed Dec 26 21:30:05 2007
Author: rana 
Type: Laser 
Category: PSL 
Subject: laser power 
I have rewired the laser power so now the Diode Box goes directly to one of the
overhead umbilicals (circuit #29) which is energized by the wall switch which
is marked as "Master Laser". Cutting off that switch is NOT the desired way to
power down the laser; use the Beckoff touchpad instead. The wall switch is the
emergency cutoff.

The HEPA fans on the laser table are both routed to power strips on the table now.

With the laser on, the beam looks well clipped. As far as I can tell this comes
from inside the box so the next thing is to do a little more re-alignment inside
the box.

Also need some more lenses for cameras and more IR cards.

And we should slice the tires of anyone who borrows equipment without leaving a note!
ELOG V3.1.3-